# Breaking the Gold Standard: Extracting Forgotten Data under Exact Unlearning in Large Language Models

### v1

@@ Abstract
Large language models are typically trained on datasets collected from the web, which may inadvertently contain harmful or sensitive personal information. To address growing privacy concerns, unlearning methods have been proposed to remove the influence of specific data from trained models. Of these, exact unlearning—which retrains the model from scratch without the target data—is widely regarded the gold standard, believed to be robust against privacy-related attacks. In this paper, we challenge this assumption by introducing a novel data extraction attack that compromises even exact unlearning. Our method leverages both the pre- and post-unlearning models: by guiding the post-unlearning model using signals from the pre-unlearning model, we uncover patterns that reflect the removed data distribution. Combining model guidance with a token filtering strategy, our attack significantly improves extraction success rates—doubling performance in some cases—across common benchmarks such as MUSE, TOFU, and WMDP. Furthermore, we demonstrate our attack’s effectiveness on a simulated medical diagnosis dataset to highlight real-world privacy risks associated with exact unlearning. In light of our findings, which suggest that unlearning may, in a contradictory way, increase the risk of privacy leakage, we advocate for evaluation of unlearning methods to consider broader threat models that account not only for post-unlearning models but also for adversarial access to prior checkpoints.
@@ 3 Threat Model
[…]
@@ 5.1 Experimental Setup
Unless otherwise noted, we set the forgetting set size to 10% of the full dataset. For our method, the guidance scale $w$ is set to $2.0$ for Phi and $1.4$ for Llama, and the constraint level $\gamma$ is set to $10^{-5}$ by default. We analyze the impact of different fine-tuning iterations and forgetting set sizes in Sec. 5.3, and investigate the effect of varying hyper-parameters on the MUSE dataset in Sec. 5.4. Further details on training and dataset preparation are provided in Appendix Sec. A, and additional experimental results on our extraction method under approximate unlearning are presented in Appendix Sec. C.
@@ 5.2 Main Comparison
To ensure fair comparison with previous work, we adopt a baseline attack that directly generates text from the given LLMs scalable before unlearning. Following prior studies tofu ; muse , we use greedy sampling by default, as it tends to exhibit higher memorization. We evaluate our method on multiple datasets (MUSE, TOFU, WMDP) using both Phi-1.5 and Llama2-7b, with 10% of the data designated as the forgetting set. As shown in Tab. 1, our method consistently achieves substantial improvements in extraction performance across all settings. Notably, the strict extraction accuracy (A-ESR($\tau=1.0$)) doubles in some cases and increases by at least 0.4$\times$ in most settings, highlighting the effectiveness of our approach. Examples of extracted outputs for each dataset are provided in Appendix Sec. D.
@@ 5.6 Possible Defense against the Attack
(d) ROUGE-L(R)
Figure 9: A-ESR1.0
As shown in Fig. 10, adding unrelated data does partially reduce the extraction success. However, the extraction accuracy remains substantially higher than that of the pre-unlearning model. Even with 10$\times$ more unrelated data than the forgetting set—more than doubling the computational cost—the Rouge-L(R) and A-ESR metrics only exhibit a slight decline. This suggests that our extraction method is primarily instance-level and does not heavily rely on the model’s overall conceptual knowledge, making it relatively insensitive to the introduction of additional unrelated data.
@@ A.2 Dataset Preparation
  • MUSE [28]: We use the MUSE-News dataset, which consists of BBC news articles collected after August 2023. The dataset is split into two disjoint subsets: $\mathcal{D}_{\text{forget}}$ and $\mathcal{D}_{\text{retain}}$, containing 0.8M and 1.6M tokens, respectively. For a $k\%$ forgetting set, we randomly select passages from $\mathcal{D}_{\text{forget}}$ until the total number of selected tokens reaches $2.4$$\times$ $k\%$. The prefix known to the attacker is the first half of each sentence.
@@ Appendix B Medical Dataset Experiment Details
Appendix C Extraction under Approximate Unlearning
We evaluate our extraction method under several approximate unlearning techniques [21, 33, 28]. Following our default setup, we fine-tune Phi-1.5 on the TOFU dataset for 3 epochs to obtain the pre-unlearning model, and experiment with a forgetting set that makes up 10% of the full dataset. For unlearning, we follow prior work [28], using a constant learning rate of $10^{-5}$ and stopping when the post-unlearning Rouge-L(R) score drops to or below that of exact unlearning. In our setting, this condition is consistently met after one epoch; further training leads to excessive utility degradation.
  • Gradient Ascent (GA) [13, 12]: Applies gradient ascent on the cross-entropy loss to suppress the likelihood of the forget set. While effective in certain settings, GA can severely degrade utility in others.
  • Negative Preference Optimization (NPO) [33]: Modifies the offline DPO objective to treat the forget set as negative preference data, encouraging low likelihood on it while remaining close to the original model.
  • Gradient Descent on the Retain Set (GD) [21]: Adds a standard cross-entropy loss on the retain set $D_{\text{retain}}$ to maintain performance on non-forgotten data.
  • KL Divergence Minimization (KL) [21]: Encourages the unlearned model’s output distribution to remain close to that of the original model on inputs from the retain set.
We follow TOFU’s default settings [21] for all approximate unlearning hyper-parameters. Following prior work [28, 21], utility is measured using the Rouge-L(R) score on the retain set.
@@ Appendix C Extraction under Approximate Unlearning
Exact Unlearning (EU) 0.437 0.566 0.643 0.005 0.070         0.120$\uparrow$71.40.567
GA 0.235 0.566 0.569 0.000 0.070        0.073$\uparrow$4.30.240
$\text{GA}_{\mbox{GD}}$ 0.437 0.566 0.587 0.010 0.070         0.090$\uparrow$28.60.516
$\text{GA}_{\mbox{KL}}$ 0.243 0.566 0.571 0.002 0.070        0.075$\uparrow$7.10.253
NPO 0.272 0.566 0.579 0.003 0.070         0.080$\uparrow$14.30.282
$\text{NPO}_{\mbox{GD}}$ 0.282 0.566 0.580 0.003 0.070         0.080$\uparrow$14.30.293
$\text{NPO}_{\mbox{KL}}$ 0.243 0.566 0.571 0.003 0.070        0.073$\uparrow$4.30.253
Appendix D Visualization
@@
Appendix E Limitations and Broader Impact

### v2

# Rethinking Exact Unlearning under Exposure: Extracting Forgotten Data under Exact Unlearning in Large Language Models

@@ Abstract
Large Language Models are typically trained on datasets collected from the web, which may inadvertently contain harmful or sensitive personal information. To address growing privacy concerns, unlearning methods have been proposed to remove the influence of specific data from trained models. Of these, exact unlearning—which retrains the model from scratch without the target data—is widely regarded the gold standard for mitigating privacy risks in deployment. In this paper, we revisit this assumption in a practical deployment setting where both the pre- and post-unlearning logits API are exposed, such as in open-weight scenarios. Targeting this setting, we introduce a novel data extraction attack that leverages signals from the pre-unlearning model to guide the post-unlearning model, uncovering patterns that reflect the removed data distribution. Combining model guidance with a token filtering strategy, our attack significantly improves extraction success rates—doubling performance in some cases—across common benchmarks such as MUSE, TOFU, and WMDP. Furthermore, we demonstrate our attack’s effectiveness on a simulated medical diagnosis dataset to highlight real-world privacy risks associated with exact unlearning. In light of our findings, which suggest that unlearning may, in a contradictory way, increase the risk of privacy leakage during real-world deployments, we advocate for evaluation of unlearning methods to consider broader threat models that account not only for post-unlearning models but also for adversarial access to prior checkpoints. Code is publicly available at: https://github.com/Nicholas0228/unlearned_data_extraction_llm.
@@ 3 Threat Model
    $\mbox{A-ESR}_{\tau}(X_{0},\widehat{X})=\frac{1}{|X_{0}|}\sum\limits_{i=1}^{|X_{0}|}\mbox{ROUGE-L(R)}(X_{0}^{(i)},{\widehat{X}}^{(i)})\geq\tau.$ (1)
@@ 5.1 Experimental Setup
Unless otherwise noted, we set the forgetting set size to 10% of the full dataset. For our method, the guidance scale $w$ is set to $2.0$ for Phi and $1.4$ for Llama, and the constraint level $\gamma$ is set to $10^{-5}$ by default. We analyze the impact of different fine-tuning iterations and forgetting set sizes in Sec. 5.3, and investigate the effect of varying hyper-parameters on the MUSE dataset in Sec. 5.4. Further details on training and dataset preparation are provided in Appendix Sec. A. We present additional results for our extraction method under LoRA fine-tuning in Appendix Sec. C, for larger LLMs in Appendix Sec. D, and for comparisons with other extraction attacks in Appendix Sec. E.
@@ 5.2 Main Comparison
[…]
@@ 5.5 Real-World Scenario Simulation: Extraction of Patient Information
(d) ROUGE-L(R)
Figure 9: A-ESR1.0
5.6 Extraction under Approximate Unlearning
We evaluate our extraction method under several approximate unlearning techniques tofu ; npo ; muse . Following our default setup, we fine-tune Phi-1.5 on the TOFU dataset for 3 epochs to obtain the pre-unlearning model, and experiment with a forgetting set that makes up 10% of the full dataset. For unlearning, we follow prior work muse , using a constant learning rate of $10^{-5}$ and stopping when the post-unlearning Rouge-L(R) score drops to or below that of exact unlearning. In our setting, this condition is consistently met after one epoch; further training leads to excessive utility degradation.
  • Gradient Ascent (GA) jang2022knowledge ; ilharco2022editing : Applies gradient ascent on the cross-entropy loss to suppress the likelihood of the forget set. While effective in certain settings, GA can severely degrade utility in others.
  • Negative Preference Optimization (NPO) npo : Modifies the offline DPO objective to treat the forget set as negative preference data, encouraging low likelihood on it while remaining close to the original model.
  • Gradient Descent on the Retain Set (GD) tofu : Adds a standard cross-entropy loss on the retain set $D_{\text{retain}}$ to maintain performance on non-forgotten data.
  • KL Divergence Minimization (KL) tofu : Encourages the unlearned model’s output distribution to remain close to that of the original model on inputs from the retain set.
We follow TOFU’s default settings tofu for all approximate unlearning hyper-parameters. Following prior work muse ; tofu , utility is measured using the Rouge-L(R) score on the retain set.
[…]
@@ 5.7 Possible Defense against the Attack
[…]
@@ 6 Conclusion and Discussion
7 Acknowledgments
Zhiwei Steven Wu was in part supported by an NSF CAREER Award #2339775 and NSF Award #2232693.
@@ A.2 Dataset Preparation
[…]
@@ Appendix B Medical Dataset Experiment Details
Appendix C Extraction Results under LoRA Fine-tuning
In practice, parameter-efficient fine-tuning methods such as LoRA [10] are increasingly popular. We therefore examine whether our attack remains effective when the pre-unlearning model is LoRA-fine-tuned on TOFU. Specifically, we provide results on Phi-1.5 fine-tuned with LoRA on TOFU for five epochs using a constant learning rate of 2e-4 and a 10% forgetting set, while keeping all other settings consistent with Sec. 5.2. Then, we employ a guidance scale w=1.6 for our extraction, increasing the extraction rate. The improvements are consistent, as shown in Tab. 4.
Table 4: Comparison of our extraction method with baselines under LoRA fine-tuning on Phi-1.5 on TOFU dataset. The improvements are consistent across metrics.
@@ Appendix C Extraction Results under LoRA Fine-tuning
Post-Unlearning Generation 0.412 0.010 0.003
Pre-Unlearning Generation 0.544 0.070 0.055
Ours 0.614 0.160 0.108
Appendix D Extraction Results on Larger LLMs
Table 5: Comparison of our extraction method with baselines under LoRA fine-tuning on Mixtral-8x7B-Instruct-v0.1 on TOFU dataset. The improvements are consistent across metrics.
[…]
Post-Unlearning Generation 0.372 0.015 0.013
Pre-Unlearning Generation 0.465 0.040 0.018
Ours 0.537 0.103 0.055
We further experiment with a larger model, Mixtral-8x7B [16], which has substantially more parameters, employs a MoE architecture, and is instruction-tuned, making it sufficiently distinct from our default models (LLaMA and Phi) to better evaluate generalization. Specifically, we report results on Mixtral-8x7B-Instruct-v0.1111https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, fine-tuned with LoRA on TOFU for two epochs using a cosine learning rate schedule with a base learning rate of 1e-4 and a 10% forgetting set (all other settings follow Sec. 5.3). For extraction, we apply a guidance scale of $w=2$. The improvements remain consistent, as shown in Tab. 5.
Appendix E Comparison with Other Extraction Attacks
ETHICIST [36] is a method for extracting memorized training data by tuning soft prompt embeddings and applying loss smoothing with calibrated confidence estimation. Therefore, we provide experiments comparing our method with ETHICIST on the TOFU dataset with Phi-1.5 under the 10% forgetting setting. Notably, ETHICIST assumes a stronger threat model that the attacker has access to part of the model’s training data. To align with this setting, we assume the attacker has access to half of the retained set in TOFU (while remaining blind to the target extraction set, i.e., the forgetting set), and use it to train the soft prompt. The hyper-parameters of ETHICIST follow the configuration provided in the original paper’s open-sourced code.
As shown in Tab. 6, experimental results show that ETHICIST alone can partially improve the extraction rate. However, since ETHICIST is designed to extract training data from a single model, it is complementary rather than conflicting with our method, which leverages the difference between pre- and post-unlearning models. In fact, combining ETHICIST with our guidance-based approach leads to even stronger extraction performance.
Table 6: Comparison of our extraction method with ETHICIST under default fine-tuning on the TOFU dataset. Our method outperforms ETHICIST by a clear margin, and combining the two further improves performance.
[…]
Pre-Unlearning Generation 0.566 0.100 0.070
ETHICIST 0.570 0.118 0.073
Ours 0.643 0.202 0.120
Ours + ETHICIST 0.652 0.213 0.153
Appendix F Visualization
@@
Appendix G Limitations and Broader Impact
@@
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction highlight a new threat: that exact unlearning—previously regarded as the gold standard for data removal—can introduce new privacy risks. This claim is the central focus of the paper and is supported by the proposed method in Sec.  4 and a set of experiments that design and evaluate a data extraction attack leveraging both pre- and post-unlearning models in Sec. 5.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: The paper discusses the limitations in Appendix Sec. G.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper includes heuristic modeling to provide intuitive guidance for the proposed method in Sec. 4, but it does not involve formal theoretical assumptions or proofs.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: The main experimental procedures are outlined in Sec. 5, with additional implementation details provided in Appendix Sec. A, ensuring reproducibility of the results relevant to the core claims.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: The code is provided via an link shown at abstract. The datasets used are primarily open-source benchmarks. The main experimental results are presented in Sec. 5, with reproduction details provided in Appendix Sec. A.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The main experimental settings are described in Sec. 5, with additional details provided in Appendix Sec. A. Hyper-parameters and their selection process are explained in Sec. 5.4, ensuring clarity on how they were chosen.
  30. Guidelines:
  31. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  32. 7.
    Experiment statistical significance
  33. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  34. Answer: [Yes]
  35. Justification: Statistical significance is addressed in Tab. 1. The variation of A-ESR across three unlearning runs is minimal (standard deviation < 0.01), so error bars are omitted for clarity. The improvements of our extraction method over the baseline are substantially larger than this variance, indicating the results are statistically robust.
  36. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  37. 8.
    Experiments compute resources
  38. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  39. Answer: [Yes]
  40. Justification: Information available in Sec. A
  41. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  42. 9.
    Code of ethics
  43. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  44. Answer: [Yes]
  45. Justification: We have reviewed the NeurIPS Code of Ethics and ensured that the research complies with all relevant ethical guidelines.
  46. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  47. 10.
    Broader impacts
  48. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  49. Answer: [Yes]
  50. Justification: The paper discusses the broader impacts in Appendix Sec. G.
  51. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  52. 11.
    Safeguards
  53. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  54. Answer: [N/A]
  55. Justification: The paper does not center on releasing any new datasets or models, and thus no additional safeguards are required.
  56. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  57. 12.
    Licenses for existing assets
  58. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  59. Answer: [Yes]
  60. Justification: We use multiple benchmark datasets, all of which are properly cited in Sec. 5, with detailed descriptions provided in Appendix Sec. A.
  61. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  62. 13.
    New assets
  63. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  64. Answer: [N/A]
  65. Justification: The paper does not release new assets.
  66. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  67. 14.
    Crowdsourcing and research with human subjects
  68. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  69. Answer: [N/A]
  70. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  71. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  72. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  73. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  74. Answer: [N/A]
  75. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  76. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  77. 16.
    Declaration of LLM usage
  78. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  79. Answer: [N/A]
  80. Justification: LLMs are used during the construction of the medical dataset, as described in Sec. 5.5, but they are not part of the core methodology or contribution of the paper.
  81. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.